The area of Kens rectangular yd is 480 sq ft. The yard is 24 ft wide. What is the length of fencing Ken will need to buy in order to fence in the garden completely o all four sides? Explain how you got your answer.
Answer by checkley79(3341)

AREA=LW
480=L*24
480=24L
L=480/24
L=20 FT. IS THE LENGTH.
PERMITER=2L+2W
P=2*20+2*24
P=40+48
P=98 FT. OF FENCING IS NEEDED.